Two friends rent an apartment for $975 per month. Since one bedroom is 60 square feet larger than the other bedroom, each person’s rent contribution is based on bedroom size. Each person agrees to pay $3.25 per square foot of bedroom area. What is the size of the larger bedroom
Answer by rfer(16322)

975/3.25=300 sq ft total
x+x+60=300
2x=240
x=120 sq ft
x+60=180 sq ft